OpenStudy (anonymous):

I WILL MEDAL AND FAN The length of a rectangle is three times its width. If the perimeter is at most 112 centimeters, what is the greatest possible value for the width? A. 22.4 cm B. 28 cm C. 14 cm D. 42 cm

OpenStudy (anonymous):

do you need help with the steps

jigglypuff314 (jigglypuff314):

so 2w + 2(3w) <_ 112 then solve for width as if it had been an equals 2w + 2(3w) = 112 simplify 2w + 6w = 112 8w = 112 so w = 112/8 which would simplify to be your answer

OpenStudy (anonymous):

ty!!!